OpenStudy (anonymous):

Solve for x: -2|x - 3| = -12 x = -3, x = 4.5 x = 3, x = 9 Medal ((&)) Fan.. Please Help Asap !! x = -3, x = 9 No solutions

OpenStudy (nurali):

-2|x - 3| = -12

OpenStudy (anonymous):

It is your last answer choice

OpenStudy (nurali):

Divide by -2

OpenStudy (nurali):

|x - 3| = 6

OpenStudy (anonymous):

Which would be no solution ?

OpenStudy (anonymous):

I met x=-3 and x=9

OpenStudy (nurali):

|x - 3| =6 x-3=6 or x-3=-6 x=9 or x=-3

OpenStudy (anonymous):

Ohh so its C ?

OpenStudy (anonymous):

Yes you can place the numerical values of x and see if it equals

OpenStudy (anonymous):

thanks to all of you

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

i tagged you in another question as well @Nurali

OpenStudy (anonymous):

Your welcome